### v1

# FIHA: Autonomous Hallucination Evaluation in Vision-Language Models with Davidson Scene Graphs

Abstract
The rapid development of Large Vision-Language Models (LVLMs) often comes with widespread hallucination issues, making cost-effective and comprehensive assessments increasingly vital. Current approaches mainly rely on costly annotations and are not comprehensive – in terms of evaluating all aspects such as relations, attributes, and dependencies between aspects. Therefore, we introduce the FIHA (autonomous Fine-graIned Hallucination evAluation evaluation in LVLMs), which could access hallucination LVLMs in the LLM-free and annotation-free way and model the dependency between different types of hallucinations. FIHA can generate Q&A pairs on any image dataset at minimal cost, enabling hallucination assessment from both image and caption. Based on this approach, we introduce a benchmark called FIHA-v1, which consists of diverse questions on various images from MSCOCO and Foggy. Furthermore, we use the Davidson Scene Graph (DSG) to organize the structure among Q&A pairs, in which we can increase the reliability of the evaluation. We evaluate representative models using FIHA-v1, highlighting their limitations and challenges. We released our code and data [^1] https://anonymous.4open.science/r/FIHA-45BB.
1 Introduction
Table 1: Comparison with other benchmarks. Dis. denotes Discriminative and Gen. denotes Generative.
Evaluation Methods Discriminative Hallucination Task Type Use DSG LLM Free Annotation Free
Object Attribute Relation Dis. Gen.
POPE [26] $\times$ $\times$ $\times$ $\times$
NOPE [31] $\times$ $\times$ $\times$ $\times$ $\times$
CIEM [15] $\times$ $\times$ $\times$ $\times$
Bingo [8] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
AMBER [37] $\times$ $\times$
HallusionBench [28] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
MHaluBench [4] $\times$ $\times$ $\times$ $\times$ $\times$
Hal-Eavl [17] $\times$ $\times$
FIHA (ours)
Large Vision-Language Models (LVLMs) such as MiniGPT-4  [45] and LLaVA  [30], which extend Large Language Models (LLMs) by incorporating visual encoders, such as CLIP [34], have shown prominent capabilities in visual understanding and generation [42]. However, LVLMs suffer from the issue of hallucination, which can lead to misinterpretation or erroneous assertions of the visual inputs, thus hindering the performance of models in multi-modal tasks [16, 21]. Specifically, the models may describe objects that do not exist in the image or incorrect object attributes and relations between objects. Generating such unreliable content will greatly reduce the model’s credibility. Therefore, it is crucial to establish a benchmark for evaluating the hallucination level of LVLMs.
Previous studies [26, 38, 37], as shown in Table 1, primarily employ a Question Generation (QG) module to create a set of validation questions and expected answers (i.e. Q&A pairs) for hallucination evaluation. These generated questions are then used to evaluate hallucinations in LVLMs. Despite the compelling success of the existing work, they still face two main challenges: (1) The existing work overlooks the dependency between different kinds of questions. For example, if the answer to “Is there a bike?” is no, dependent questions like “Is the bike yellow?” should be skipped. (2) Additionally, most prior work heavily relies on human annotations [37] or LLMs [24] to generate Q&A pairs used in hallucination evaluation, which can be costly or labor-intensive.
To mitigate these limitations, we propose Fine-grained Hallucination Evaluation (FIHA), an automatic evaluation framework for assessing fine-grained and diverse hallucinations in large-scale vision-language models. The framework accepts either images or captions as input and generates Q&A pairs by extracting objects, attributes, and entity relations. It then formulates diverse questions (e.g., “what”, “who”, “which”, etc.) that allow for free-form responses. By integrating BLIP-2 [25] for caption generation, Fast R-CNN [11] for feature extraction, and a question-generation template, our pipeline enables fully automatic Q&A generation without relying on LLMs or manual annotations.
To organize all the Q&A pairs into a tree-like structure, we introduce the Davidson Scene Graph (DSG) [5]. The DSG ensures that responses to leaf node questions depend on the correctness of the root node answers. This structure mitigates situations where a large language model (LLM) may provide correct answers to fine-grained questions while hallucinating basic attributes. By enforcing this dependency, we aim to reduce errors and improve the reliability of the model’s responses. In addition, our Q&A pairs cover various types of questions, including negative, narrative, and interrogative questions. This type of tree structure allows for a progressive deepening of questions, enabling a comprehensive evaluation of the model’s understanding of the image.
We make the following key contributions through this work:
  • To the best of our knowledge, FIHA is the first automated hallucination evaluation framework that is LLM-free and annotation-free. This approach not only scales efficiently but also minimizes labor and associated costs.
  • Based on FIHA, we generate a DSG-based fine-grained evaluation benchmark FIHA-v1 that includes Q&A pairs evaluating various types of hallucinations and the semantic dependency relation organized by DSG.
  • We evaluate and analyze several mainstream open-source and close-source LVLMs with FIHA-v1, providing valuable insights into their performance.
2 Methodology
Figure 1: Overview of FIHA framework. FIHA extracts entities, attributes, and relations from images and captions respectively, and generates comprehensive and diverse questions to thoroughly detect model hallucinations. In the Figure, we can see that no large language model (LLM) [1] or additional manual annotations are used.
In this section, we introduce the overall pipeline of FIHA as illustrated in Figure 1. In summary, our pipeline offers two approaches for Q&A generation. The first is based on images: given an image $I$, we extract the necessary entities, including features such as objects, object attributes, and relations. Using a rule-based method, we then generate the Q&A pairs. The second approach is based on captions. If an image does not already have a caption, we can use BLIP-2 [25] to generate captions. Alternatively, if the dataset includes original captions, we can use them directly as input, pass them through the feature extraction process, and generate the corresponding Q&A pairs.
2.1 Fine-grained Information Extraction
2.1.1 Information Acquisition from the Caption
Q&A generation based on caption includes caption generation (optional if original datasets include captions) and extract information (object existence, object attributes and object relations) and using these information to generate Q&A pairs.
Caption Generation. Image captions can depict an image in detail, demonstrating fine-grained visual information, such as objects, attributes and relations. Inspired by the findings of previous research [26], which indicate that smaller vision-language models tend to produce more concise responses with fewer hallucinations compared to mainstream LVLMs. As such, we select BLIP-2 to generate a caption for the image. This step allows us to generate highly credible captions based on the image.
Fine-grained Information Extraction. In this process, we take either the generated captions (if the ground-truth caption is not available) or the ground-truth captions, depending on the user’s needs, as input and extract information such as object existence, object attributes, and relations from the captions. For extracting objects and attributes in the caption, we use SpaCy’s [14] part-of-speech tagging feature to identify objects and their corresponding attributes, such as numerals, adjectives, and verbs. As a result, we obtain all the ground truth objects and their attributes as: $G^{C}_{O,A}=\{o_{1}:A_{1},o_{2}:A_{2},\dots,o_{n}:A_{n}\}$, where $n$ is the number of objects. $o_{i}$ is the $i$-th object and $A_{i}$ is all attributes for the $i$-th object. Relations from the captions are extracted using the Stanford CoreNLP library [^2] https://stanfordnlp.github.io/CoreNLP/., which provides a powerful suite of NLP tools for performing various linguistic analyses on text, making it an ideal choice for relation extraction. From this process, we obtain all the relations: $G^{C}_{R}=\{R_{1}(o_{R_{1}}^{1},o_{R_{1}}^{2}),R_{2}(o_{R_{2}}^{1},o_{R_{2}}^{2}),\dots,R_{m}((o_{R_{m}}^{1},o_{R_{m}}^{2}))\}$, where $m$ is the number of relations. $R_{i}$ is the $i$-th relation for the objects $o_{R_{i}}^{2}$ and $o_{R_{i}}^{2}$.
2.1.2 Information Acquisition from the Image
As the caption may lose some information in the image, our second approach to extract information is directly from images, which provides richer and more detailed information than captions alone. For object and attribute extraction, we use Fast R-CNN [11], a well-established and widely used object detection method based on Region-based Convolutional Networks. Fast R-CNN has been a pioneering approach in the field of object detection due to its ability to quickly identify objects within an image while simultaneously predicting their attributes. This method allows us to retrieve the ground truth objects along with their corresponding attributes such as color, size, and shape, forming a set of objects and attributes: $G^{I}_{O,A}=\{o_{1}:A_{1},o_{2}:A_{2},\dots,o_{n}:A_{n}\}$, where $n$ represents the number of objects detected. In addition to identifying objects and their attributes, it is crucial to understand how these objects interact or relate within the scene. For this purpose, we employ RelTR [6], a cutting-edge method designed to generate sparse scene graphs by decoding visual appearances and learning both subject and object queries from the image data. RelTR enables us to extract meaningful relationships between the detected objects, such as spatial relations (e.g., one object being behind or near another) and actions (e.g., wearing or holding), resulting in a set of ground truth relations: $G^{I}_{R}=\{R_{1}(o_{R_{1}}^{1},o_{R_{1}}^{2}),R_{2}(o_{R_{2}}^{1},o_{R_{2}}^{2}),\dots,R_{m}((o_{R_{m}}^{1},o_{R_{m}}^{2}))\}$, where $m$ denotes the number of relationships extracted from the image. By combining both object detection and relational extraction, this approach provides a comprehensive understanding of the visual content, which is essential for generating accurate and meaningful Q&A pairs.
Figure 2: Example of extracted information.
2.2 Question Answer Pair Generation
We use two kinds of questions for hallucination evaluation. The first type is Yes-No questions, which check object existence, such as “Is there any {objk}?” and object relations like “Is there a {obj1} near the {objk}?”. These questions help determine whether specific objects and their relationships are present within an image. Additionally, Yes-No questions assess attributes by asking about features like color, size, or location.
The second type is Wh-Questions, which add diversity to the evaluation by incorporating interrogative words such as “what”, “who”, “which”, “where”, and “how many”. These questions elicit more detailed, free-form responses, typically no longer than three words. For example, “What color is the {objk}?” or “Which object is near the {objk}?” help assess the finer details about objects and their relationships. Unlike traditional hallucination evaluations that primarily rely on Yes-No questions, our approach includes both types to provide a more comprehensive assessment.
Additionally, we introduce Negative Questions for both Yes-No questions and Wh-Questions. These questions are generated by replacing the objects, their attributes, and relations in the original Q&A pairs with non-existent ones. Specifically, we randomly select several pieces of information from a complete list of objects, attributes (such as color, size, etc.), and relations to replace the real information. Such questions are answered with negative pronouns such as “none”, “nobody”, or “nowhere”.
2.3 Davidson Scene Graph
To model the dependency between objects, attributions, and relations accurately and improve the reliability of hallucination evaluation, we introduce the Davidsonian Scene Graph (DSG) [5] mechanism. The DSG can be understood as a post-processing step for the Q&A pairs. After obtaining all the Q&A pairs, we organize them into multiple tree-like structures, where each Q&A pair serves as a node. According to the structure of the tree-like structures, each node is either a root node or a leaf node. Specifically, the entire process is divided into three steps. In step 1, we set the question about the existence of a certain object as the root node. In step 2, we set all questions related to the object of the root node, such as those about its attributes and relations, as corresponding leaf nodes. Finally, in step 3, determine whether the root node question is answered correctly; if not, there is no need to judge the questions at the leaf nodes, and we directly determine that all questions on the tree are answered incorrectly. For instance, after step 1 and step 2, we obtain a list of questions such as $L^{Q}=\{Q_{1}:Independent,Q_{2}:Depends\ on\ Q_{1}\}$. Before determining if the answer to $Q_{2}$ is correct, we first assess $Q_{1}$, which concerns the accuracy related to the root node. If the question about the existence of an object, which is at the root node, is answered incorrectly, we consider that all other related questions must be hallucinatory.
3 Experiments
Table 2: Data statistics of FIHA-v1. #Num denotes the number of the images.
Source #Num From Image From Caption
MSCOCO 500 25,699 13,007
Foggy 150 7,232 2,801
Visual Genome 50 1,566 476
Table 3: Evaluation results of LVLMs on questions generated from images and captions using FIHA. The upper part is from the MSCOCO dataset and the bottom part is from the Foggy dataset.
Model Question Generated from Image Question Generated from Caption
Acc. P. R. F1 Acc. P. R. F1
MSCOCO
mPLUG-Owl 42.1 70.2 61.4 43.7 31.4 61.6 55.5 31.2
MiniGPT-4 23.5 27.5 22.2 22.1 15.9 25.7 28.8 14.2
MultiModal-GPT 59.1 46.4 47.1 46.6 23.8 39.6 45.7 22.1
LLaVA-1.5-7B 77.8 77.0 65.9 67.7 50.7 64.9 67.5 50.5
LLaVA-1.5-13B 78.9 80.9 66.4 68.3 47.6 64.2 65.5 48.5
InstructBLIP 84.7 83.3 78.6 80.4 65.7 69.5 77.4 64.2
GPT-4V 87.2 81.4 86.3 85.5 70.3 71.5 75.8 69.3
Foggy
mPLUG-Owl 64.8 60.2 51.1 42.7 29.5 58.9 51.6 25.6
MiniGPT-4 30.1 30.2 27.6 28.1 23.4 34.4 37.8 23.0
MultiModal-GPT 50.2 48.7 46.1 45.8 28.1 43.9 47.9 25.4
LLaVA-1.5-7B 67.7 68.4 56.2 52.9 29.1 50.0 49.2 25.8
LLaVA-1.5-13B 68.1 71.5 56.1 52.3 28.9 49.2 49.8 25.5
InstructBLIP 70.9 75.6 60.2 58.8 32.8 58.3 53.2 30.5
GPT-4V 76.3 70.1 64.6 66.0 33.7 53.3 51.7 32.1
Visual Genome
mPLUG-Owl 41.8 68.9 60.9 43.3 44.6 71.7 51.8 33.8
MiniGPT-4 22.9 26.8 22.0 21.8 15.9 25.7 28.8 14.2
MultiModal-GPT 58.8 46.1 46.9 46.3 65.3 65.2 62.2 61.7
LLaVA-1.5-7B 77.7 77.2 61.1 67.9 56.4 73.8 62.0 52.6
LLaVA-1.5-13B 79.0 81.2 66.7 68.6 74.3 79.6 68.8 69.2
InstructBLIP 84.5 83.7 79.0 80.7 67.7 78.4 71.9 66.7
GPT-4V 87.0 81.2 86.0 85.3 84.2 78.9 84.1 82.2
3.1 Setup
Datasets. We construct a hallucination evaluation benchmark FIHA-v1 based on three datasets: the MSCOCO [27], the Foggy [7] and Visual Genome [22]. MSCOCO is a large image dataset developed by Microsoft, officially known as Microsoft Common Objects in Context. This dataset aims to advance the development of computer vision tasks such as object detection, segmentation, and image captioning. This dataset contains over 330,000 images, of which more than 200,000 images are annotated, covering 80 different object categories. Foggy is a synthetic fog dataset that simulates fog in real-world scenes. Each foggy image is rendered using clear images and depth maps from Cityscapes. There are a total of 1,500 images and 500 for each three different foggy levels of each cityscape image: no foggy, medium foggy, and dense foggy. Visual Genome [22] is a large image datasets which includes 108,077 images. Those images have some overlap with the MSCOCO.
Metrics. We use Accuracy (Acc.), Precision (P.), Recall (R.), and F1 Score (F1) as evaluation metrics for Yes-No questions. For Wh-Questions, we use the Accuracy, Recall, and F1 Score from BERTScore [43] for evaluation.
Models We select seven mainstream LVLMs for evaluation: mPLUG-Owl [40], MiniGPT-4 [45], MultiModal-GPT [12], LLaVA-1.5-7B [30], LLaVA-1.5-13B [30], InstructBLIP [10], and GPT-4V [32].
3.2 Data Processing and Analysis
We randomly selected 500 images from the MSCOCO dataset, 150 images from the Foggy, and 50 from the Visual Genome dataset. Using the process described in Section 2, we generate tens of thousands of Q&A pairs. The detailed data statistics of our FIHA-v1 benchmark can be found in Table 2.
Figure 3: Distribution of two types of question, i.e. Yes-No and Wh-questions.
Figure 3 illustrates the distribution of question types generated from images and captions. The proportion of questions related to objects, attributes, and relations is relatively balanced, reflecting the rationality of the method design. It is noteworthy that the abundance of the question category reflects FIHA’s effective capability in generating tasks of the generation type, thereby enabling a more effective assessment of hallucinations.
Table 4: The results of a more fine-grained assessment of LVLMs from the perspectives of object, attribute, and relation. The results are based on statistics from Q&A pairs generated from captions in the MSCOCO and Foggy Cityscapes dataset.
Model Object Attribute Relation
Acc. P. R. F1 Acc. P. R. F1 Acc. P. R. F1
MSCOCO
mPLUG-Owl 57.3 75.7 47.3 48.0 20.6 55.7 53.5 20.4 22.7 56.5 55.8 22.7
MiniGPT-4 66.2 59.5 62.6 59.5 9.6 12.8 9.2 9.4 4.7 12.1 11.4 4.9
MultiModal-GPT 51.6 54.1 51.5 42.5 16.0 39.2 42.8 15.8 12.1 30.8 39.6 11.8
LLaVA-1.5-7B 79.2 82.4 77.5 78.4 27.9 55.6 56.7 27.8 47.9 59.1 69.7 44.7
LLaVA-1.5-13B 70.8 80.6 70.2 68.3 34.3 56.4 59.7 33.7 42.1 58.3 66.6 48.1
InstructBLIP 84.6 87.7 81.4 84.2 61.0 62.2 76.2 55.6 57.5 61.0 75.7 52.1
GPT-4V 90.8 87.7 89.8 88.6 83.6 77.7 85.2 79.8 66.2 61.2 73.2 58.3
Foggy
mPLUG-Owl 52.9 32.3 50.0 39.2 15.7 54.8 52.1 15.3 11.8 34.6 46.9 11.1
MiniGPT-4 62.1 60.6 58.4 57.8 9.6 25.1 14.6 9.3 8.5 23.2 26.5 8.5
MultiModal-GPT 52.9 59.7 52.6 42.1 12.6 33.9 38.6 12.5 11.5 33.3 39.4 11.4
LLaVA-1.5-7B 54.0 63.3 54.0 44.4 11.5 33.2 46.0 10.8 15.4 47.8 48.9 15.1
LLaVA-1.5-13B 54.2 62.8 54.2 44.6 11.3 31.4 46.3 10.6 14.9 47.0 48.6 14.6
InstructBLIP 54.2 65.2 53.9 44.2 20.7 55.1 54.6 20.6 15.9 48.5 49.2 15.6
GPT-4V 61.8 69.6 59.2 54.5 11.1 37.0 33.1 11.0 20.4 50.5 50.4 20.3
3.3 Results
3.3.1 Overall Results on Datasets Generated by FIHA
We show the hallucination comparison of the seven mainstream LVLMs on our FIHA-v1 in Table 3. From this Table, we have several observations. 1) It’s worth highlighting that GPT-4V excels in both image and caption Q&A pairs, achieving the best performance among the evaluated models. 2) The second-best performer is InstructBLIP, which significantly outperforms other models except GPT-4V across most metrics. 3) Additionally, we have observed that model parameters are also significant factors affecting performance. For instance, LLaVA-1.5-13B provides a more comprehensive improvement over the LLaVA-1.5-7B.
In addition, we also show the performance of 7 mainstream LVLMs on FIHA-v1 based on the Visual Genome dataset. The results show a similar trend as compared to the performance in MSCOCO datasets. Specifically, the GPT-4V performs best and MiniGPT-4 performs the worst. LLaVA-1.5-13B performs better than LLaVA-1.5-7B, which also indicates that the model parameter size influences the performance.
3.3.2 Fine-Grained Results
Furthermore, we evaluate the model’s performance from more dimensions (i.e. the object existence, attribute, and relation) with FIHA. We show the fine-grained evaluation results in Table 4.
Object Hallucination
From the results for the object, we can observe that even after introducing more negative samples, the Accuracy and Precision of the models remain high, indicating that most models have a strong capability to determine whether an object exists or not. In comparison, the Recall is somewhat lower, indicating that the model still has a tendency to lean towards affirmative responses.
Attribute Hallucination
From the results for the attribute, tt is evident that this part of the hallucination is much more difficult to identify. Compared to the object itself, its color, quantity, size, and so on are indeed more challenging to judge. Even the best-performing GPT-4V has an F1 score of less than 80 on regular data. Moreover, the performance of the vast majority of models plummets on special datasets, indicating that the robustness of existing LVLMs needs to be enhanced.
Relation Hallucination
From the results for the relation, this part is the most challenging, with the F1 score of GPT-4V on regular data not even reaching 60%. The potential reason is that Q&A pairs of the relation types involved more than one object, which makes it challenging.
4 Analysis
In this section, we further evaluate the effectiveness of our benchmark FIHA-v1 by four research questions.
4.1 How Reliable is FIHA?
To assess the reliability of the benchmark, we manually evaluate the accuracy of the Q&A pairs in the FIHA-v1 generated by the pipeline. This evaluation involves verifying the correctness of the Q&A pairs by checking whether the answers match the corresponding questions. We show the evaluation results in Table 5, showing that the Q&A pairs generated from image captions are highly reliable, achieving a 96% accuracy rate in samples from the MSCOCO dataset. The pipeline based on MSCOCO, which uses Fast R-CNN, achieves a 98.2% accuracy rate for Q&A pairs. However, the remaining 1.8% error can lead to incorrect Q&A pairs. For example, the model sometimes fails to identify specific details, such as ears in an image. Another challenge relates to feature extraction. For instance, in one case, a horse’s color is incorrectly identified as black when it is actually white or light gray. Although such inaccuracies are rare, they highlight a weakness of the framework and underscore the need for ongoing advancements in object detection and feature extraction technologies. Overall, the FIHA framework demonstrates high reliability in generating datasets for evaluating hallucinations in LVLMs, with the dataset generated from captions performing exceptionally well and achieving near-perfect accuracy.
Table 5: The accuracy of Q&A pairs generated from different datasets.
Source From Image From Caption
MSCOCO 98.2 96.0
No Foggy 98.1 96.1
Medium Foggy 97.6 94.5
Dense Foggy 96.3 94.1
In addition, we conduct tests on more complicated images. Specifically, we use the Foggy dataset [7], which contains three-level foggy for an image: no foggy, medium foggy, and dense foggy. We show the cases in this dataset in Appendix B. We select a total of 150 images, with 50 images for each of the three foggy levels. This comparison aims to assess the influence of noise on the accuracy of the framework. For the Q&A pair generation from images approach, as shown in Table 5, in the dense fog condition, the accuracy for Q&A pairs generation from images approach is as high as 96.3%. For Q&A pairs generation from captions approach, the accuracy is 94.1%. For medium fog conditions, it is 97.6% and 94.5%; and for no fog conditions, the accuracy are 98.1% and 94.1%. The no-fog result show a similar accuracy with MSCOCO. From the results, we can see that as the degree of fog increases, the accuracy of FIHA’s Q&A pairs decreases, indicating the challenge posed by blurry images. Notably, despite the image with dense fog, Q&A pairs based on our FIHA still has a high accuracy rate.
4.2 What is the Impact of Introducing the DSG?
To reasonably improve the reliability of hallucination assessment, we introduce the DSG mechanism, which can model the dependency between different types of hallucination. As introduced in the previous section, by reorganizing the problem into a tree structure, the judgment of each leaf node depends on the correctness of the root node’s judgment. In this section, we quantitatively analyze the impact brought by the DSG.
Table 6 shows the performance difference for each model before and after introducing DSG. It shows that stronger models like GPT-4V and LLaVA are less affected, while the metrics for other models have dropped by more than 50%. The reason might be that these weaker models make a lot of mistakes in the root-level questions which mostly ask about the object’s existence and answer the leaf-level questions correctly by random guessing. Therefore, after the introduction of the DSG, they are marked as failing on all related leaf node questions, leading to a significant drop in performance. This indicates the reliability of introducing the DSG mechanism.
Table 6: The performance decrease of various LVLMs after introducing DSG.
Model Acc.$\downarrow$ P.$\downarrow$ R.$\downarrow$ F1$\downarrow$
mPLUG-Owl 29.6 22.1 14.0 28.7
MiniGPT-4 62.6 51.8 62.1 61.2
MultiModal-GPT 21.3 27.6 21.9 24.3
LLaVA-1.5-7B 4.2 11.7 4.5 4.8
LLaVA-1.5-13B 2.7 8.1 3.3 3.6
InstructBLIP 5.7 9.6 5.7 5.7
GPT-4V 6.0 9.9 5.4 8.4
4.3 Is the Information Extracted from Images More Comprehensive?
As shown in Figure 1, we extract information from both the image and the caption to construct Q&A pairs. Typically, the image itself contains more abundant information. In this section, we will verify whether the information extracted from the image is more comprehensive and diverse than that extracted from the caption. We have separately counted the number of six different types of Q&A pairs from image and caption, mainly focusing on the three directions of object, attribute, and relation. As shown in Figure 4, it is evident that the information extracted from the image surpasses the information extracted from the caption.
Figure 4: Comparison of the number of Q&A paris across different types of hallucination from image and caption.
4.4 Why are Performance on our Benchmark Lower Than Others?
In the experiment, we observe that our test results are lower than others, e.g. POPE [26] and HaELM [38], indicating that FIHA can detect more difficult and distinct issues. We analyze that there are mainly three reasons: firstly, we added a large number of misleading negative samples, and since the model tends to give affirmative answers (Section 3.3.1), this increases the difficulty of evaluation. Secondly, the role of DSG directly impacts the results and improves the reliability of the evaluation method. Finally, the comprehensiveness of FIHA is more challenging than methods that focus primarily on generating coarse-grained object-level questions.
5 Related Work
In this section, we mainly discuss existing Large Vision-Language Models (LVLMs) and the hallucination problems that exist in LVLMs.
5.1 Large Vision-Language Model
With the success of pretraining techniques in Large Language Models (LLMs) [36] and Vision Foundation Models (VFMs) [3], many researchers [2, 23] expanded LLMs to comprehend real-world images through LVLMs with advance in-context or few-shot learning capabilities. As a result, there has been a surge in visual instruction-adapted LVLMs [30, 45, 10, 12], demonstrating remarkable generalization performance across various Vision-Language (VL) tasks. Most of these studies utilized GPT-4 to generate multimodal instruction tuning datasets and multi-stage pretraining to align the visual information with the pre-trained LLM. For example, Liu et al. [30] utilized the visual encoder output as input for LLaMA [36] and trained both networks to align on the generated visual instruction dataset. Zhu et al. [45] integrated Vicuna [33] as a language decoder and only fine-tuned the cross-modal alignment network with extended image captions from ChatGPT. Likewise, both Multimodal GPT [12] and InstructBLIP [10] used various instruction-tailored VL datasets. In addition, the former adopted BLIP2 [25] as its foundational architecture while the latter initialized from Flamingo [2].
Despite the advancements of LVLMs, they remain encumbered by the persistent challenge of hallucinations when generating textual output. These issues significantly hinder their effectiveness in various vision-language tasks [35].
5.2 Hallucination in LVLMs
Recently, there has been growing research attention directed towards the phenomenon of hallucination in LVLMs. Among these works, some studies, as shown in Table 1, have concentrated on hallucination detection and evaluation [26, 38, 37, 20], and some have developed methods to mitigate hallucination [29, 44, 41, 19]. For example, the Bingo [9] benchmark evaluates hallucinations in GPT-4V by focusing on bias and interference, such as regional bias and vulnerability to leading questions. HallusionBench [13] provides a comprehensive diagnostic suite for evaluating entangled language hallucination and visual illusions, emphasizing visual commonsense and reasoning with diverse visual-question control pairs. AutoHallusion [39] introduces an automatic method for generating hallucination benchmarks by manipulating images to conflict with language priors, evaluating models on both synthetic and real-world data. Hal-Eval [18] offers a universal framework categorizing hallucinations into objects, attributes, relations, and events, utilizing a detailed taxonomy to assess LVLMs. Though the issue of hallucination is studied extensively, only a few works have focused on fine-grained hallucination detection in LVLMs. For instance, Li et al. [26] proposed a novel evaluation metric “POPE” to evaluate hallucinations in LVLMs by pooling questions about the ground truth objects. They showed that existing state-of-the-art LVLMs are highly prone to object-level hallucinations. Wang et al. [38] introduced “HaELM”, a framework for detecting hallucinations. They utilized LLMs to generate a hallucinatory dataset and then fine-tuned LLaMA to identify hallucinatory responses from LVLMs. The aforementioned line of research either exclusively focused on object-level hallucination or required training for the detection of hallucination. To address these challenges, Wang et al. [37] introduced “AMBER”, a comprehensive benchmark capable of assessing both generative and discriminative tasks based on object attribute and relation hallucination. Though this work developed a fine-grained hallucination evaluation framework, it required human annotators to annotate the object existence, object attribute, and object relation information for discriminative tasks.
In contrast to the aforementioned studies, our work differs by being applicable to any existing dataset or unseen images for generating probing questions related to object existence, attributes, and relations for evaluating LVLM hallucination. Instead of relying on human annotators, we use an object detection model that performs better in object detection tasks than LLMs. In addition, we consider the dependent relation between object, attribute, and relation.
6 Conclusion
In recent years, large vision-language models have developed quickly, but hallucinations remain a serious concern. Current hallucination evaluation methods face problems like high costs, limited scope, and lack of generalization. Thus, we introduce FIHA, a multi-dimensional detection method that requires no LLMs and no annotations. FIHA can automatically create high-quality Q&A pairs for any image dataset. We conducted a thorough analysis of the performance of mainstream LVLMs, identified the issues, and proposed potential methods for improvement. In the future, we will delve deeper into methods for alleviating hallucinations.
Appendix A Code Example for Generating QA Pairs Based on Extracted Information
1if relation.endswith(tuple([’ing’, ’ed’])):
2 question = f"Isthe{subject}{relation}the{object}intheimage?"
3elif relation.endswith(tuple([’over’, ’under’, ’above’, ’near’, ’behind’, ’on’, ’at’])):
4 if obj_is_living(object):
5 question = f"Whois{relation}the{object}intheimage?"
6 else:
7 question = f"Whatis{relation}the{subject}intheimage?"
Appendix B Example of foggy Cityscapes Images datasets
Figure 5: no foggy
Figure 6: medium foggy
Figure 7: dense foggy